# Camera-based 3D Semantic Scene Completion with Sparse Guidance Network

### v1

@@ Introduction
By learning features with hybrid guidance, we obtain discriminative voxel features with rich semantic context and spatial geometry cues. Then we design the multi-scale semantic diffusion (MSSD) module to propagate the semantic information from seed features to the whole scene based on the geometry and spatial occupancy cues. The MSSD module contains three anisotropic convolution layers Li et al. (2020) and the ASPP module, which is lightweight while well capturing multi-scale features for instances with different sizes. Afterward, we use the head consisting of a linear layer and softmax layer to predict the final semantic scene prediction $\hat{\textbf{Y}}\in\mathbb{R}^{C_{class}\times X\times Y\times Z}$ from the diffused voxel features.
@@
Following MonoScene Cao and de Charette (2022), we adopt the Scene-Class Affinity Loss on both semantic and geometry results of $\hat{\textbf{Y}}$ to optimize the class-wise derivable precision, recall, and specificity metrics simultaneously. The overall loss function for the final prediction is formulated by:
@@
Training Loss
Unlike VoxFormer Li et al. (2023d) with sophisticated two-stage training, we train our SGN end-to-end. The total training loss $\mathcal{L}=\mathcal{L}_{geo}+\mathcal{L}_{occ}+\mathcal{L}_{sem}+\mathcal{L}_{ssc}$.
Experiments
To evaluate our SGN, we conduct extensive experiments on the challenging large-scale outdoor dataset SemanticKITTI Behley et al. (2019). Due to the page limitation, more details on the datasets, metrics, experiments, and qualitative results are provided in the supplementary materials.
Implementation Details
We crop the input RGB images of cam2 to size $1220\times 370$ and take the 2D feature map with 1/16 of the input resolution. The feature dimension $C$ and the channel number $C_{o}$ are set to 128 and 8, respectively. The size $X\times Y\times Z$ of the 3D feature volume is $128\times 128\times 16$. And the final predictions are up-sampled to $256\times 256\times 32$. The threshold $\theta$ for selecting seed voxels is set to 0.5 by default. We train SGN for 40 epochs on 4 V100 GPUs with a total batch size of 4. The AdamW Loshchilov and Hutter (2017) optimizer is used with an initial learning rate of 2e-4 and a weight decay of 1e-2. Following VoxFormer Li et al. (2023d), we design the single-image version SGN-S, taking only the current image as input and the temporal version SGN-T with the current and the previous 4 images as input. We also provide the light temporal version SGN-L, which takes temporal inputs and uses ResNet18 as the backbone with dimension $C=64$ and 1 anisotropic convolution layer for MSSD.
Variants MSSD SG GG VA IoU (%) mIoU (%)
@@
Table 3: Ablation on network components
@@
Table 4: Impact of different training modes. Memory denotes training memory
@@
Table 5: Ablation on image features
Module IoU (%) mIoU (%) Params (M) Memory (G)
FLoSP 43.60 14.55 28.16 14.21
LSS 42.95 14.77 30.56 13.23
Cross-Attention 43.66 14.05 62.15 19.01
Table 6: Ablation on view transformation. Memory denotes training memory
Comparison with the state-of-the-art
@@
Table 1 and Table 2 present the comparison results between our SGN and other state-of-the-art camera-based SSC methods on the SemanticKITTI validation and test sets, respectively. Our SGN-T achieves state-of-the-art performance on both SemanticKITTI validation and test sets. Specifically, SGN-T outperforms the second one by 1.86% points (OccFormer) and 2.19% points (VoxFormer) regarding mIoU and IoU, as shown in Table 2. And compared with these fully dense processing methods with complex 3D models, such as MonoScene, OccFormer, and Symphonize, our SGN-S also performs better in terms of mIoU and IoU. For example, SGN-S greatly boosts the MonoScene by 3.47% points in mIoU and 6.74% points in IoU, demonstrating the effectiveness of our dense-sparse-dense design equipped with hybrid guidance. Notably, SGN-S outperforms the recent VoxFormer by 2.2% points in mIoU but has a slightly lower IoU (-0.42% points). We explain that VoxFormer adopted a two-stage training approach, and the first stage was trained offline with temporal inputs, helping enhance occupancy precision. However, our SGN-S is end-to-end trained with only a single frame as input. Compared to the two-stage VoxFormer, the higher mIoU score of our end-to-end SGN-S for semantic scene completion demonstrates the superiority of our framework of semantic propagation based on geometry prior and occupancy information.
@@
Remarkably, our light version SGN-L achieves notable performance (45.45 IoU and 15.32 mIoU) on SemanticKITTI validation with only 12.5M parameters. Compared with MonoScene, OccFormer, and VoxFormer with $\sim$150M, $\sim$200M, and $\sim$60M parameters, our SGN-L performs better while being more lightweight. It demonstrates that our SGN requires no heavy 3D model and has a more powerful representation ability.
Ablation studies
We do ablation studies on network components, training mode, image features, view transformation, and seed voxels on SemanticKITTI validation. All experiments are conducted with our SGN-S by default.
Ablation on network components.
We do ablation studies to analyze the effect of the proposed semantic guidance (SG), geometry guidance (GG), multi-scale semantic diffusion (MSSD), and voxel aggregation layer (VA) in Table 3. GP and OA denote geometry prior and occupancy-aware information, respectively. Firstly, we construct a baseline that directly attaches a segmentation head after the selected seed features. As shown in the first line of Table 3, the constructed baseline has already achieved 41.76% IoU and 10.62% mIoU scores, demonstrating our depth-based sparse voxel proposal network can provide effective seed voxels. When equipped with our MSSD (Variant 1), the IoU and mIoU scores are improved by 1.46% points and 2.38% points, respectively. It demonstrates the effectiveness of multi-scale information propagation. And our semantic guidance on the seed features further boosts the mIoU by 0.68% points (Variant 2 vs. Variant 1), showing the importance of intra-category separation of seed features. On the other hand, the geometry guidance brings slight improvement in terms of IoU score, while the geometry prior further boosts the mIoU score by 0.95% points (Variant 4 vs. Variant 3). Besides, introducing occupancy-aware information can help improve performance (Variant 5 vs Variant 4). And comparing Variant 5 with Variant 1, the mIoU score is significantly boosted (+1.55% points), demonstrating the effectiveness of our semantic diffusion based on geometry prior and occupancy information.
Impact of different training modes.
To explore the effect of different training modes, i.e., two-stage and one-stage training, we provide the detailed comparison results with VoxFormer in Table 4. Line 3 of Table 4 presents the results of our SGN-S with the two-stage training strategy. Note that when equipping SGN-S with a two-stage approach, the first stage remains the same as VoxFormer, and the parameters and training memory of the second stage are calculated for a fair comparison. In the same training configuration, Our SGN-S surpasses VoxFormer-S by a large margin regarding mIoU scores (+2.58% points). Even our one-stage SGN-S outperforms two-stage VoxFormer-S by 2.2% points in mIoU. It is worth noting that the model parameters of our SGN-S are only about half of those of VoxFormer-S. For the temporal version, our SGN-T achieves 46.21 IoU and 15.32 mIoU scores, boosting VoxFormer-T by 2.06% points in IoU and 1.97% points in mIoU. Our light version, SGN-L, with only 12.5M parameters, also outperforms VoxFormer-T on mIoU and IoU scores while requiring only about half the training memory (7.16 G).
Impact of image features.
The 2D features provide a foundation for informative voxel features. We do detailed experiments in Table 5 to explore the impact of the feature scale and image backbone. We see that using 2D features at a scale of 1/16 in ResNet 50 achieves the best IoU score and has comparable mIoU to other variants. And it strikes a balance between performance and model size. The results of using ResNet18 as the backbone are presented in the last line of Table 5 and shows that a lighter image backbone reduces the model parameters by 12.43M while the performance of the mIoU score drops by 0.47% points.
Ablation on view transformation.
The view transformation generates the initial 3D features for the subsequent hybrid guidance and informative voxel features. We further investigate the effect of different view transformation modules. We implement three commonly used modules, i.e., FLoSP in Monoscene, LSS Philion and Fidler (2020), and cross-attention adopted from VoxFormer. The results are shown in Table 6, showing that using FLoSP contains fewer parameters while achieving comparable performance to other variants on both mIoU and IoU scores. Therefore, our SGN lifts the 2D features to 3D volume with the view transformation designed in the spirit of FLoSP.
Figure 4: Impact of threshold values for seed voxels.
Exploration on the threshold for seed voxels.
We change the value $\theta$ to investigate the impact of different thresholds for selecting seed voxels. We calculate the average occupancy rate of seed voxels, mIoU score, and IoU score for variants with $\theta$ from 0.1 to 0.9. The results are shown in Figure 4, showing that the performance of the model first increases and then decreases as $\theta$ increases. And when $\theta=0.4$, the model achieves the best performance on both mIoU and IoU scores. Interestingly, we found our model still achieves notable mIoU and IoU scores when the occupancy ratio of the seed voxels is very low ($<5\%$ points). It shows that seed voxels with high confidence play an essential role in our semantic diffusion.
Conclusion
This work focuses on camera-based semantic scene completion (SSC). Existing methods usually rely on sophisticated 3D models to directly process the coarse lifted 3D features that are not discriminative enough for clear segmentation boundaries. Therefore, we propose the end-to-end SGN to diffuse semantics from the semantic- and occupancy-aware seed voxels to the whole scene based on geometry prior and occupancy information. By designing hybrid guidance and effective voxel aggregation, we enhance the intra-category feature separations and expedite the convergence of semantic diffusion. Extensive experiments on the SemanticKITTI benchmark demonstrate the effectiveness of Our SGN achieves state-of-the-art performance while being more lightweight.
Appendix A Appendix
Dataset and Metrics
Dataset.
For large-scale outdoor scene understanding, the KITTI odometry dataset Geiger, Lenz, and Urtasun (2012) collects 22 sequences with 20 classes with a Velodyne HDL-64 laser scanner in the scenes of autonomous driving. SemanticKITTI Behley et al. (2019) is based on the KITTI dataset and provides semantic annotation of all sequences. According to the official setting for semantic scene completion (SSC), sequences 00-07 and 09-10 (a total of 3834 scans) are for training, sequence 08 (815 scans) is for validation, and the rest (3901 scans) is for testing. The volume of interest for the SSC benchmark is $[0\sim 51.2m,-25.6m\sim 25.6m,-2\sim 4.4m]$, and the voxelization resolution s is $0.2m$. The SSC labels with resolution $256\times 256\times 32$ of train and validation set are provided for the users. In this work, we focus on the camera-based SSC, taking the RGB images as inputs similar to Cao and de Charette (2022); Li et al. (2023d); Jiang et al. (2023).
Metrics.
Following Song et al. (2017), we mainly report the Intersection-over-Union (IoU) for scene completion and mIoU of $C_{n}=19$ classes (no “unlabeled” class) for semantic scene completion. The mIoU is calculated by:
$mIoU=\frac{1}{C_{n}}\sum_{c=1}^{C_{n}}{\frac{TP_{c}}{TN_{c}+FP_{c}+FN_{c}}}$ (10)
Methods Modality IoU (%) mIoU (%)
12.8m 25.6m 51.2m 12.8m 25.6m 51.2m
SSCNet Song et al. (2017) LiDAR 64.37 61.02 50.22 20.02 19.68 16.35
JS3CNet Yan et al. (2021) LiDAR 63.47 63.40 53.09 30.55 28.12 22.67
MonoScene* Cao and de Charette (2022) Camera 38.42 38.55 36.80 12.25 12.22 11.30
OccFormer Zhang, Zhu, and Du (2023) Camera 56.38 47.28 36.50 20.91 17.90 13.46
VoxFormer-T Li et al. (2023d) Camera 65.38 57.69 44.15 21.55 18.42 13.35
SGN-S (ours) Camera 64.21 56.20 43.60 21.53 19.60 14.55
SGN-L (ours) Camera 70.08 61.17 45.45 24.76 21.17 14.80
SGN-T (ours) Camera 70.61 61.90 46.21 25.70 22.02 15.32
Table 7: Quantitative comparison in different ranges. “*” denotes the results provided by VoxFormer.
Further Analysis
We conduct a series of experiments on SemanticKITTI validation to provide further analysis on ranges, model dimensions, and temporal input. The efficiency analysis is also provided.
Quantitative comparison in different ranges.
@@ Appendix A Appendix
We also provide the results of different ranges in Table 7. The results show that SGN-T achieves mIoU scores of 25.70% and 22.02% within 12.8 meters and 25.6 meters and performs better than VoxFormer-T by 4.15% and 3.60% points in mIoU, respectively. Additionally, our SGN-S surpassed MonoScene by 9.28% and 7.38% points in mIoU within 12.8 meters and 25.6 meters. Notably, SGN-T obtains competitive performance with LiDAR-based methods in short-range (12.8 meters) areas. For example, SGN-T outperforms SSCNet by 5.58% points in mIoU and 6.24% points in IoU within 12.8 meters, demonstrating the effectiveness of our camera-based method for autonomous driving.
Number of model dimensions.
The impact of the number of model dimensions of our SGN-S is evaluated and presented in Table 8. The results reveal that using a large number of feature channels for 3D features boosts the models’ performance while increasing the model complexity. For example, the model with 64 dimensions contains fewer parameters and requires less training memory, although its performance drops by 0.28% points in IoU and 0.35% points in mIoU (Line3 vs Line4). We also provide the results of different depths for the multi-scale semantic diffusion module. We see that using different depths has comparable mIoU and IoU, demonstrating that our model with hybrid guidance and semantic diffusion avoids the dependency on the heavy 3D model for processing 3D features.
Effect of temporal input.
Finally, we explore the impact of the number of temporal frames in Table 9. We take historical frames to form the temporal input. As Table 9 shows, the model’s performance on IoU scores first increases with the number of frames and then decreases. We explain that the camera extrinsic matrix from the historical frame to the current system is not always accurate, especially when the temporal interval is long. Therefore, the 3D points may project on the wrong image patches of the historical frames, which disturbs the learning of voxel features.

### v2

@@ III-D Multi-Scale Semantic Propagation
By learning features with hybrid guidance and voxel aggregation, we obtain discriminative voxel features $\textbf{F}_{f}^{3D}$ with the rich semantic context in the seed features $\textbf{F}_{s}^{3D}$ and spatial geometry cues from previous 3D features $\textbf{F}_{n}^{3D}$, and occupancy-aware features $\textbf{F}_{o}^{3D}$. Then we design the multi-scale semantic propagation (MSSP) module to propagate the semantic information from seed features to the whole scene. The MSSP module contains three anisotropic convolutional layers [13] and the ASPP [57] module consisting of three 3D convolutions with the kernel size of $3\times 3\times 3$ and dilation of 1, 2, and 4. This module is lightweight and can well capture multi-scale features of instances of different sizes through convolutional kernels with different receptive fields. Afterward, we use the head consisting of a linear layer and softmax layer to predict the final semantic scene prediction $\hat{\textbf{Y}}\in\mathbb{R}^{C_{class}\times X\times Y\times Z}$ from the propagated voxel features.
@@
Following MonoScene [7], we adopt the Scene-Class Affinity Loss to force the network to account for voxels within the same category as well as voxels across different categories. The Affinity Loss optimizes the class-wise derivable precision, recall, and specificity metrics simultaneously, where precision and recall evaluate the performance of voxels within the same class, while specificity assesses the performance of dissimilar voxels. Specifically, similar to [7, 32], we apply scene-class affinity loss on both semantic and geometry results of the prediction $\hat{\textbf{Y}}$. We integrate and optimize scene- and class-wise semantics $\mathcal{L}_{scal}^{sem}$, geometry $\mathcal{L}_{scal}^{geo}$, and cross-entropy loss $\mathcal{L}_{ce}$. The overall loss function is formulated by:
@@
III-E Training Loss
Unlike VoxFormer [12] with sophisticated two-stage training, we train our SGN end-to-end. The total training loss $\mathcal{L}=\mathcal{L}_{geo}+\mathcal{L}_{occ}+\mathcal{L}_{sem}+\mathcal{L}_{ssc}$.
IV Experiments
In this section, we present the datasets, evaluation metrics, and detailed implementation aspects of our approach. Subsequently, we conduct extensive experiments to establish that our proposed SGN consistently surpasses or achieves comparable performance against the state-of-the-art methods on the complex, large-scale outdoor dataset SemanticKITTI [14] as well as SSCBench-KITTI-360 [15]. Following this, we provide qualitative results to underscore the efficacy of our SGN. Moreover, we conducted detailed ablation studies to dissect the contribution of individual components of our method and various configurations, thereby offering an in-depth analysis of our approach. Additionally, we provide detailed experiments on the NYUv2 dataset [58] to demonstrate the generalization ability of our SGN on indoor scenes.
IV-A Dataset and Metrics
Dataset
For large-scale outdoor scene understanding, the KITTI odometry dataset [50] collects 22 sequences with 20 classes with a Velodyne HDL-64 laser scanner in the scenes of autonomous driving. SemanticKITTI [14] is based on the KITTI dataset and provides semantic annotation of all sequences. According to the official setting for semantic scene completion (SSC), sequences 00-07 and 09-10 (a total of 3834 scans) are for training, sequence 08 (815 scans) is for validation, and the rest (3901 scans) is for testing. SSCBench-KITTI-360 [15] offers a comprehensive benchmark for semantic scene completion, featuring nine densely annotated sequences of urban driving scenes. The dataset is meticulously partitioned, with the training set encompassing 8,487 frames across scenes 00, 02-05, 07, and 10. The validation set is carefully curated with 1,812 frames from scene 06, ensuring a robust evaluation framework. Furthermore, the testing set includes 2,566 frames exclusively from scene 09, providing a diverse and challenging environment for model assessment. The volume of interest for both two SSC benchmarks is $[0\sim 51.2m,-25.6m\sim 25.6m,-2\sim 4.4m]$, and the voxelization resolution s is $0.2m$. The SSC labels with resolution $256\times 256\times 32$ of train and validation set are provided for the users. In this work, we focus on the camera-based SSC, taking the RGB images as inputs similar to [7, 12, 30].
Metrics
Following [16], we mainly report the Intersection-over-Union (IoU) for scene completion and mIoU of $C_{n}$ classes (no “unlabeled” class) for semantic scene completion. The mIoU is calculated by:
$mIoU=\frac{1}{C_{n}}\sum_{c=1}^{C_{n}}{\frac{TP_{c}}{TN_{c}+FP_{c}+FN_{c}}}$ (13)
IV-B Implementation Details
We crop the input RGB images of cam2 to size $1220\times 370$ for SemanticKITTI and RGB images of cam1 of $1408\times 376$ for SSCBench-KITTI-360. The 2D feature maps with 1/16 of the input resolution are taken for the subsequent processing. The feature dimension $C$ and the channel number $C_{o}$ are set to 128 and 8, respectively. The size $X\times Y\times Z$ of the 3D feature volume is $128\times 128\times 16$. And the final predictions are up-sampled to $256\times 256\times 32$. The threshold $\theta$ for selecting seed voxels is set to 0.5 by default. We train SGN for 40 epochs on 4 V100 GPUs with a total batch size of 4. The AdamW [59] optimizer is used with an initial learning rate of 2e-4 and a weight decay of 1e-2. Following VoxFormer [12], we design the single-image version SGN-S, taking only the current image as input and the temporal version SGN-T with the current and the previous 4 images as input. We also provide the lightweight version SGN-L, which takes temporal inputs and uses ResNet18 as the backbone with dimension $C=64$ and 1 anisotropic convolution layer for MSSP.
TABLE V: Ablation on network components, i.e., semantic guidance (SG), geometry guidance (GG), multi-scale semantic propagation (MSSP), and voxel aggregation layer (VA).
Variants MSSP SG GG VA IoU (%) mIoU (%)
@@ IV-B Implementation Details
TABLE VI: Impact of different training modes. Our one-stage SGN surpasses the two-stage VoxFormer by a large margin. Memory denotes training memory.
@@
TABLE VII: Ablation study for depth estimator. Mono and Stereo denote using monocular-based Adabins [60] and stereo-based MobileStereoNet [53] as the depth estimator.
Methods Depth IoU (%) mIoU (%)
12.8m 25.6m 51.2m 12.8m 25.6m 51.2m
VoxFormer-S Mono 57.41 50.61 38.68 14.62 14.01 10.67
Stereo 65.35 57.54 44.02 17.66 16.48 12.35
VoxFormer-T Mono 59.03 50.47 38.08 18.67 15.42 11.27
Stereo 65.38 57.69 44.15 21.55 18.42 13.35
SGN-S (ours) Mono 59.82 51.43 39.35 18.13 16.65 12.51
Stereo 64.21 56.20 43.60 21.53 19.60 14.55
SGN-L (ours) Mono 63.79 54.99 41.36 20.77 17.69 12.64
Stereo 70.08 61.17 45.45 24.76 21.17 14.80
SGN-T (ours) Mono 64.74 55.55 41.87 21.43 17.94 12.91
Stereo 70.61 61.90 46.21 25.70 22.02 15.32
TABLE VIII: Ablation on view transformation. Using FLoSP contains fewer parameters while achieving comparable performance to other variants. Memory denotes training memory.
Module IoU (%) mIoU (%) Params (M) Memory (G)
FLoSP 43.60 14.55 28.16 14.21
LSS 42.95 14.77 30.56 13.23
Cross-Attention 43.66 14.05 62.15 19.01
TABLE IX: Ablation on image features, including the image backbone and scales. Using scale 16 strikes a balance between performance and model parameters.
@@
IV-C Comparison with the state-of-the-art
SemanticKITTI
@@ IV-C Comparison with the state-of-the-art
Table I and Table II present the comparison results between our SGN and other state-of-the-art camera-based SSC methods on the SemanticKITTI validation and test sets, respectively. Our SGN-T achieves state-of-the-art performance on both SemanticKITTI validation and test sets. Specifically, SGN-T outperforms the second one by 1.86% points (OccFormer) and 2.19% points (VoxFormer) regarding mIoU and IoU, as shown in Table II. And compared with these fully dense processing methods with complex 3D models, such as MonoScene and OccFormer, our SGN-S also performs better in terms of mIoU and IoU. For example, SGN-S greatly boosts the MonoScene by 3.05% points in mIoU and 6.48% points in IoU, demonstrating the effectiveness of our dense-sparse-dense design equipped with hybrid guidance. Notably, SGN-S outperforms the recent VoxFormer by 2.2% points in mIoU but has a slightly lower IoU (-0.42% points). We explain that VoxFormer adopted a two-stage training approach, and the first stage was trained offline with temporal inputs, helping enhance occupancy precision. However, our SGN-S is end-to-end trained with only a single frame as input. Compared to the two-stage VoxFormer, the higher mIoU score of our one-stage SGN-S for semantic scene completion demonstrates the superiority of our framework of semantic propagation based on spatial geometry cues.
@@
Remarkably, our lightweight version SGN-L achieves notable performance (45.45% IoU and 14.80% mIoU) on SemanticKITTI validation with only 12.5M parameters. Compared with MonoScene, OccFormer, and VoxFormer with $\sim$150M, $\sim$200M, and $\sim$60M parameters, our SGN-L performs better while being more lightweight. It demonstrates that our SGN requires no heavy 3D model and has a more powerful representation ability.
@@
Quantitative comparison in different ranges. We also provide the results of different ranges in Table III. The results show that SGN-T achieves mIoU scores of 25.70% and 22.02% within 12.8 meters and 25.6 meters and performs better than VoxFormer-T by 4.15% and 3.60% points in mIoU, respectively. Additionally, our SGN-S surpassed MonoScene by 9.28% and 7.38% points in mIoU within 12.8 meters and 25.6 meters. Notably, SGN-T obtains competitive performance with LiDAR-based methods in short-range (12.8 meters) areas. For example, SGN-T outperforms SSCNet by 5.68% points in mIoU and 6.24% points in IoU within 12.8 meters, demonstrating the potential application of our camera-based method for autonomous driving.
SSCBench-KITTI-360
Table IV presents the comparison results between our SGN and other state-of-the-art SSC methods, including LiDAR-based methods (SSCNet, LMSCNet) and camera-based methods (MonoScene, TPVFormer, VoxFormer, OccFormer, DepthSSC, Symphonize) on the SSCBench-KITTI-360 test sets. We can see that our SGN outperforms most camera-based methods by a large margin in terms of both mIou and IoU metrics. For example, SGN-S, SGN-L, and SGN-T surpass OccFormer by 5.95%, 6.37%, 6.79% points in IoU and 3.9%, 3.14% 4.44% points in mIoU. Compared to the LiDAR-based methods, our SGN-T also archives comparable performance in IoU and performs better in mIoU, demonstrating the superiority of our SGN. Interestingly, we found that our SGN-T achieves better performance on many thing classes such as traffic-sign, other-object, trucks, bicycles, motorcycles, and other vehicles while performing worse on plain stuff classes such as road, parking, sidewalk, building, and vegetation than LiDAR-based method SSCNet. We explain that the LiDAR point cloud contains more accurate structure information, which may facilitate the occupancy prediction of plain classes, while the vision feature includes more semantic information that helps distinguish objects that belong to different classes.
Figure 4: Visual comparison of our SGN-T with state-of-the-art methods on SemanticKITTI validation. Compared to VoxFormer-T and MonoScene, our SGN-T generates more precise segmentation boundaries (labeled in red circles).
IV-D Qualitative Visualizations
We provide the visualization results of the proposed SGN-T on SemanticKITTI validation in Figure 4. Compared to VoxFormer-T and MonoScene, our SGN-T generates more precise segmentation boundaries, especially on “plane” classes and large objects such as cars. Besides, SGN-T predicts more accurate SSC results and preserves more regional details in the short-range areas than other methods. For example, there are some wrong semantics and missing objects for VoxFormer-T and MonoScene in the short-range areas, as shown in the first and third rows in Figure 4. However, we noticed that our SGN-T also missed some distant objects that are very small in RGB images. We explain that our SGN uses image features of the 1/16 scale, which may degrade the performance of objects in distant areas. We also provide qualitative results of our SGN-S, SGN-L, and SGN-T on SemanticKITTI hidden test set in Figure 6. Our method can provide accurate semantic occupancy prediction and road layout and handle typical driving scenes such as crowded cars, shadows, tiny poles, and crossroads.
IV-E Ablation studies
We do ablation studies on network components, training mode, depth estimator, image features, view transformation, seed voxels, model dimensions, and temporal input on SemanticKITTI validation. All experiments are conducted with our SGN-S by default.
Ablation on network components
We do ablation studies to analyze the effect of the proposed semantic guidance (SG), geometry guidance (GG), multi-scale semantic propagation (MSSP), and voxel aggregation layer (VA) in Table V. GP and OA denote geometry information from 3D features $\textbf{F}^{3D}$ and occupancy-aware features $\textbf{F}_{o}^{3D}$, respectively. Firstly, we construct a baseline that directly attaches a segmentation head after the selected seed features. As shown in the first line of Table V, the constructed baseline has already achieved 41.76% IoU and 10.62% mIoU scores, demonstrating our depth-based sparse voxel proposal network can provide effective seed voxels. When equipped with our MSSP (Variant 1), the IoU and mIoU scores are improved by 1.46% points and 2.38% points, respectively. It demonstrates the effectiveness of multi-scale information propagation. And our semantic guidance on the seed features further boosts the mIoU by 0.68% points (Variant 2 vs. Variant 1), showing the importance of intra-category separation of seed features. On the other hand, the geometry guidance brings slight improvement in terms of IoU score, while the geometry prior further boosts the mIoU score by 0.95% points (Variant 4 vs. Variant 3). Besides, introducing geometry information in features $F_{o}^{3D}$ can help improve performance (Variant 5 vs Variant 4). And comparing Variant 5 with Variant 1, the mIoU score is significantly boosted (+1.55% points), demonstrating the effectiveness of our semantic propagation based on spatial geometry cues.
TABLE X: Number of model dimensions and depth of MSSP. Memory denotes training memory.